OpenStudy (anonymous):

five less than four times a number is the same as nine times the number. What is The Number

OpenStudy (anonymous):

let n = the number. 4n-5 = n+4 3n=9 n=?

OpenStudy (anonymous):

so the answer is ?

OpenStudy (isaiah.feynman):

@drehenderson solve it to know!

OpenStudy (anonymous):

9

OpenStudy (isaiah.feynman):